1.       Case Corporation produces cellular phone cases. Each case requires a keypad which it also manufactures at a cost per P20 per unit, inclusive of fixed overhead costs of P5.

 

Case Corporation needs 50,000 units of this keypad annually. A supplier, Keypad Corp., has offered to sell to Case Corp. Its keypad requirements at P24 per unit. If Case decides to buy the keypads, P2 per unit of the fixed overhead based on the annual estimate could be eliminated, and the facility previously used to produce the keypad could be rented to another company.

 

a.       If Case Corp. outsource the keypads but does not rent the unused facility, it would        .


1
Expert's answer
2021-06-17T09:48:25-0400

Solution:

a.). If Case Corp. outsources the keypads but does not rent the unused facility, it would incur an annual loss of P250,000 and a loss per unit of P5.


This is calculated as follows:

Total fixed overhead costs = 50000 units "\times" 5 = 250,000

Avoidable fixed cost if outsourced = 50000 units "\times" 2 = 100,000

Unavoidable fixed costs = 250,000 – 100,000 = 150,000

 

Total cost to buy = (50000 "\times" 24) + 150,000 = 1,200,000 + 150,000 = 1,350,000

Total cost to make = 50000 units "\times" 20 = 1,000,000

Total loss annually if outsourced = 1,350,000 – 1,000,000 = P250,000


Loss per unit = "\frac{250,000}{50,000} = P5" per unit